OpenStudy (anonymous):

A ladder that is 17 feet long is 8 feet from the base of a wall. How far up the wall does the ladder reach?

OpenStudy (anonymous):

15... use pythagorean theorum

OpenStudy (anonymous):

oh shoot.. i lied

OpenStudy (anonymous):

it's 5 radical 5

OpenStudy (anonymous):

Isn't it 15? 289-64=x^2=225=15